OpenStudy (anonymous):

The area of a regular octagon is 35 cm2. What is the area of a regular octagon with sides five times as long? A. 625 cm2 B. 875 cm2 C. 175 cm2 D. 245 cm2

OpenStudy (lilg132):

if 35 is the area we want to work out first what each side is because an octogan has eight sides we divide this by 8 35/8 = 4.375 is the length of the sides now it says that the length of the sides is 5 times as long so we multiply this side by 5 to make it 5 times as long 4.375 * 5 = 21.875 now we need to work out the area so we multiply it by 8 because an octogan has 8 sides 21.875 * 8 = 175 cm2

OpenStudy (lilg132):

dont forget to click best response :)

OpenStudy (anonymous):

Thank you

OpenStudy (lilg132):

your welcome